### v3

# Entropic (Gromov) Wasserstein Flow Matching with GENOT

Abstract
Optimal transport (OT) theory has reshaped the field of generative modeling: Combined with neural networks, recent Neural OT (N-OT) solvers use OT as an inductive bias, to focus on “thrifty” mappings that minimize average displacement costs. This core principle has fueled the successful application of N-OT solvers to high-stakes scientific challenges, notably single-cell genomics. N-OT solvers are, however, increasingly confronted with practical challenges: while most N-OT solvers can handle squared-Euclidean costs, they must be repurposed to handle more general costs; their reliance on deterministic Monge maps as well as mass conservation constraints can easily go awry in the presence of outliers; mapping points across heterogeneous spaces is out of their reach. While each of these challenges has been explored independently, we propose a new framework that can handle, natively, all of these needs. The generative entropic neural OT (GENOT) framework models the conditional distribution $\pi_{\varepsilon}(\mathbf{y}|\mathbf{x})$ of an optimal entropic coupling $\pi_{\varepsilon}$, using conditional flow matching. GENOT is generative, and can transport points across spaces, guided by sample-based, unbalanced solutions to the Gromov-Wasserstein problem, that can use any cost. We showcase our approach on both synthetic and single-cell datasets, using GENOT to model cell development, predict cellular responses, and translate between data modalities.
Machine Learning, Optimal Transport
1 Introduction
Mapping a probability distribution onto another is a ubiquitous challenge in machine learning, with many implications in generative modeling. Optimal transport (OT) has arisen in a few years as a major purveyor of tools to better address these challenges, both in theory and practice. OT focuses on finding maps that can effectively transform a distribution of matter onto another, by minimizing a certain notion of cost (Santambrogio, 2015). Applying OT to large-dimensional problems arising in machine learning and sciences has necessitated various modifications and adaptations. Starting with solvers that can solve approximate matching problems at large scales (Cuturi, 2013; Peyré et al., 2016; Scetbon et al., 2021, 2022), a recent plethora of OT-inspired training approaches for neural networks have been proposed (Makkuva et al., 2020; Korotin et al., 2020; Asadulaev et al., 2022; Fan et al., 2020; Uscidda & Cuturi, 2023; Lipman et al., 2023; Tong et al., 2020, 2023b). Mirroring these new possibilities in the field of single-cell genomics, arguably one of the most exciting transfers of OT to science, the use of OT has evolved from point-wise matching problems (Schiebinger et al., 2019; Demetci et al., 2022), to neural-based approaches that can, for instance, predict the response of cells to various perturbations (Bunne et al., 2021, 2022). As a testament of their success, practitioners demand ever more flexibility from Neural-OT solvers, in a suite of increasingly complex challenges:
From Deterministic to Stochastic Maps. The classic OT pipeline rests on deterministic (Monge) maps. Such maps lack flexibility, both at estimation and inference time. Practitioners may favor, instead, stochasticity, which would account naturally for instance, for the non-determinism of cell evolutions (Elowitz et al., 2002). Stochastic formulations can also produce a conditional distribution that can be used to quantify uncertainty. In the discrete setting, this property is fulfilled by entropy-regularized OT (EOT) (Cuturi, 2013). In a neural setting, some works (Korotin et al., 2022, 2023b; Asadulaev et al., 2022) have explored computing weak OT coupling by solving a min-max problem.
From Linear to Quadratic Problems. Classic “linear” OT puts in correspondence distributions defined in the same space $\mathcal{X}$, using a cost $c(\mathbf{x},\mathbf{y})$ for any two points $\mathbf{x},\mathbf{y}\in\mathcal{X}$ (most often set to be the squared Euclidean distance (Santambrogio, 2015, §1.3) in applications). Yet, more challenging use-cases arising in practice now involve source and target distributions that do not live in the same space, e.g. $\mathcal{X}$ and $\mathcal{Y}$ have differing dimensions, as in Demetci et al. (2022). The challenge, in that case, lies in the absence of a “natural” cost function between vectors of different sizes. This can be bypassed using quadratic losses (Mémoli, 2011; Sturm, 2020), a.k.a. Gromov-Wasserstein (GW) problems. Despite the challenges inherent to solving GW, practitioners have expressed much interest in that flexibility, as evidenced by the wide adoption of the mixed linear/quadratic fused Gromov-Wasserstein (FGW) distance (Vayer et al., 2018), see e.g. Klein et al. (2023); Lange et al. (2023); Nitzan et al. (2019); Zeira et al. (2022). To our knowledge, the only neural formulation for GW proposed thus far by (Nekrashevich et al., 2023) learns deterministic maps for the inner product costs, using a min-max-min optimization procedure.
Flexibility in Mass Conservation. In many real-world applications, the data acquisition process can be error-prone, resulting in outliers. To mitigate this, unbalanced OT (UOT) formulations that can discard observations have been proposed (Frogner et al., 2015; Chizat et al., 2018; Séjourné et al., 2021), with numerous applications to generative modeling (Balaji et al., 2020; Yang & Uhler, 2019; Eyring et al., 2023) and single-cell genomics (Schiebinger et al., 2019; Eyring et al., 2022; Lübeck et al., 2022).
Contributions. We propose a flexible neural OT framework that satisfies all the requirements above:
  • We propose GENOT, the first method that parameterizes linear and quadratic EOT couplings by modeling their conditional distributions with neural networks (Prop. 3.1), using flow matching (Lipman et al., 2023) as a core block, and any cost function between samples.
  • By showing that an unbalanced EOT problem is equivalent to a balanced one between re-weighted measures (Prop. 3.2), that can be estimated consistently (Prop. 3.3), we introduce U-GENOT to solve unbalanced problems.
  • We extend (U-)GENOT to solve the fused problem (§ 3.3). To our knowledge, GENOT is the first method to solve a continuous fused problem.
  • We show that GENOT outperforms other linear neural EOT solvers on the Gushchin et al. (2023) benchmark,
  • We showcase GENOT’s ability to handle crucial single-cell biology challenges: we (i) quantify lineage branching events in the developing mouse pancreas, (ii) predict cellular responses to drug perturbations, along with a well-calibrated uncertainty estimation, and (iii) introduce a novel method to translate ATAC-seq to RNA-seq data.
2 Background
Notations. We consider in this work two compact sets $\mathcal{X}\subset\mathbb{R}^{p}$, $\mathcal{Y}\subset\mathbb{R}^{q}$, referred to as the source and the target domain, respectively. In general,  $p\neq q$. The set of positive (resp. probability) measures on $\mathcal{X}$ is denoted by $\mathcal{M}^{+}(\mathcal{X})$ (resp. $\mathcal{M}_{1}^{+}(\mathcal{X})$). For $\pi\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})$, $\pi_{1}:=p_{1}\sharp\pi$ and $\pi_{2}:=p_{2}\sharp\pi$ denote its marginals. Then, for $\mu\in\mathcal{M}^{+}(\mathcal{X}),\nu\in\mathcal{M}^{+}(\mathcal{Y})$,  $\Pi(\mu,\nu)=\{\pi:\,\pi_{1}=\mu,\,\pi_{2}=\nu\}$. Moreover, $\tfrac{\mathop{}\!\mathrm{d}\mu}{\mathop{}\!\mathrm{d}\nu}$ denotes the relative density of $\mu$ w.r.t. $\nu$, and we write $\mu=\tfrac{\mathop{}\!\mathrm{d}\mu}{\mathop{}\!\mathrm{d}\nu}\cdot\nu$. For $\rho,\gamma\in\mathcal{M}^{+}(\mathcal{X})$, $\mathrm{KL}(\rho|\gamma)=\int_{\mathcal{X}}\log(\frac{\mathop{}\!\mathrm{d}\rho}{\mathop{}\!\mathrm{d}\gamma})\mathop{}\!\mathrm{d}\rho-\int_{\mathcal{X}}\mathop{}\!\mathrm{d}\gamma+\int_{\mathcal{X}}\mathop{}\!\mathrm{d}\rho$.
Linear Entropic OT. Let $c:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ be a cost function, $\mu\in\mathcal{M}_{1}^{+}(\mathcal{X}),\nu\in\mathcal{M}_{1}^{+}(\mathcal{Y})$ and $\varepsilon\geq 0$. The linear entropy-regularized OT problem reads
$\min_{\pi\in\Pi(\mu,\nu)}\int_{\mathcal{X}\times\mathcal{Y}}c\mathop{}\!\mathrm{d}\pi+\varepsilon\mathrm{KL}(\pi|\mu\otimes\nu)\,.$ (LEOT)
(LEOT) is also known as the static Schrödinger bridge (SB) problem (Léonard, 2013). With $\varepsilon=0$, we recover the Kantorovich (1942) problem. When $\mu$ and $\nu$ are discrete, we can solve (LEOT) with the Sinkhorn algorithm (Cuturi, 2013). For measures on $n$ points, its complexity is $\mathcal{O}(n^{2})$ in time, and $\mathcal{O}(n)$ or $\mathcal{O}(n^{2})$ in memory, depending on $c$.
Quadratic Entropic OT. As opposed to considering an inter-domain cost defined on $\mathcal{X}\times\mathcal{Y}$, quadratic entropic OT is concerned with seeking couplings based on intra-domain cost functions $c_{\mathcal{X}}:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ and $c_{\mathcal{Y}}:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}$:
$\min_{\pi\in\Pi(\mu,\nu)}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}$ $Q_{c_{\mathcal{X}},c_{\mathcal{Y}}}\mathop{}\!\mathrm{d}(\pi\otimes\pi)+\varepsilon\,\text{KL}(\pi\|\mu\otimes\nu),$ (QEOT)
where $Q_{c_{\mathcal{X}},c_{\mathcal{Y}}}(\mathbf{x},\mathbf{y},\mathbf{x}^{\prime},\mathbf{y}^{\prime}):=|c_{X}(\mathbf{x},\mathbf{x}^{\prime})-c_{Y}(\mathbf{y},\mathbf{y^{\prime}})|^{2}$ measures the pointwise cost distortion. With $\varepsilon=0$, we recover the Gromov-Wasserstein (Mémoli, 2011) problem. This is the standard OT formulation for measures supported on incomparable spaces. Using $\varepsilon>0$ offers computational advantages, since for discrete $\mu$, $\nu$, we can solve  (QEOT) with a mirror-descent scheme iterating the Sinkhorn algorithm (Peyré et al., 2016). This solver has $\mathcal{O}(n^{2})$ or $\mathcal{O}(n^{3})$ time complexity, depending on $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (Scetbon et al., 2022, Alg. 1 & Alg. 2), and $\mathcal{O}(n^{2})$ memory complexity.
Unbalanced Extensions. The EOT formulations presented above can only handle measures with the same total mass. Unbalanced optimal transport (UOT) (Liero et al., 2018; Chizat et al., 2018) lifts this constraint by penalizing the deviation of $p_{1}\sharp\pi$ to $\mu$ and $p_{2}\sharp\pi$ to $\nu$ with a divergence. Using the $\mathrm{KL}$ and introducing weightings $\lambda_{1},\lambda_{2}>0$ the unbalanced extension of (LEOT) reads
$\min_{\pi\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})}$ $\int_{\mathcal{X}\times\mathcal{Y}}c\mathop{}\!\mathrm{d}\pi+\varepsilon\mathrm{KL}(\pi|\mu\otimes\nu)$ (ULEOT)
$+\lambda_{1}\mathrm{KL}(\pi_{1}|\mu)+\lambda_{2}\mathrm{KL}(\pi_{2}|\nu).$
This problem can be solved efficiently in the discrete setting using a variant of the Sinkhorn algorithm (Frogner et al., 2015; Séjourné et al., 2023a). Analogously, quadratic OT also admits an unbalanced generalization, which reads
$\begin{split}\min_{\pi\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})}&\int_{(\mathcal{X}\times\mathcal{Y})^{2}}Q_{c_{\mathcal{X}},c_{\mathcal{Y}}}\mathop{}\!\mathrm{d}(\pi\otimes\pi)\,+\varepsilon\mathrm{KL}^{\otimes}(\pi|\mu\otimes\nu)\\ &+\lambda_{1}\mathrm{KL}^{\otimes}(\pi_{1}|\mu)+\lambda_{2}\mathrm{KL}^{\otimes}(\pi_{2}|\nu),\end{split}$ (UQEOT)
where $\mathrm{KL}^{\otimes}(\rho|\gamma)=\mathrm{KL}(\rho\otimes\rho|\gamma\otimes\gamma)$. This can be solved using an extension of Peyré et al. (2016)’s scheme introduced by Séjourné et al. (2023b). Each solver has the same time and memory complexity as its balanced counterpart. For both (ULEOT) and (UQEOT), instead of selecting $\lambda_{i}$, we introduce $\tau_{i}=\tfrac{\lambda_{i}}{\lambda_{i}+\varepsilon}$ s.t. we recover the hard marginal constraint for $\tau_{i}=1$, when $\lambda_{i}\rightarrow+\infty$. We write $\tau=(\tau_{1},\tau_{2})$.
Flow Matching. Given a prior $\rho_{0}\in\mathcal{M}_{1}^{+}(\mathbb{R}^{d})$ and a time-dependent vector field $(v_{t})_{t\in[0,1]}$, one can define a probability path $p_{t}$ starting from $\rho_{0}$ using the flow $\phi_{t}$ solving
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\phi_{t}(\mathbf{z})=v_{t}(\phi_{t}(\mathbf{z})),\quad\phi_{0}(\mathbf{z})=\mathbf{z},$ (1)
by setting $p_{t}=\phi_{t}\sharp\rho_{0}$. We then say that $v_{t}$ generates the path $p_{t}$. Continuous Normalizing Flows (Chen et al., 2018) (CNFs) model $v_{t,\theta}$ with a neural network , which is trained to match a terminal condition $p_{1}=\rho_{1}\in\mathcal{M}_{1}^{+}(\mathbb{R}^{d})$. Conditional Flow Matching (CFM) (Lipman et al., 2023) is a simulation-free technique to train CNFs by constructing individual paths between samples, and minimizing
$\mathbb{E}_{t,Z_{0}\sim\rho_{0},Z_{1}\sim\rho_{1}}[\|v_{t,\theta}\left([Z_{0},Z_{1}]_{t}\right)-(Z_{1}-Z_{0})\|_{2}^{2}],$ (2)
where $[Z_{0},Z_{1}]_{t}:=(1-t)Z_{0}+tZ_{1}$ with $t\sim\mathcal{U}([0,1])$. If this loss is 0, the flow maps $\rho_{0}$ to $\rho_{1}$, i.e., $\phi_{1}\sharp\rho_{0}=\rho_{1}$.
3 Generative Entropic Neural OT
In this section, we introduce GENOT, a method to learn EOT couplings by learning their conditional distributions. In § 3.1, we focus on balanced OT, and show that GENOT approximates linear or quadratic EOT couplings, solutions to (LEOT) or (QEOT) respectively. Second, in § 3.2, we extend GENOT to the unbalanced setting by loosening the conservation of mass constraint. We define U-GENOT, which approximates solutions to (ULEOT) and (UQEOT). Finally, in § 3.3, we highlight that GENOT also adresses a fused problem, combining (LEOT) and (QEOT).
3.1 Learning Entropic Couplings with GENOT
Let $\mu\in\mathcal{M}_{1}^{+}(\mathcal{X})$, $\nu\in\mathcal{M}_{1}^{+}(\mathcal{Y})$ and $\pi^{\star}_{\varepsilon}$ be an EOT coupling between $\mu$ and $\nu$, which can be a solution of problem (LEOT) or (QEOT). By the measure disintegration theorem:
$\mathop{}\!\mathrm{d}\pi^{\star}_{\varepsilon}(\mathbf{x},\mathbf{y})=\mathop{}\!\mathrm{d}\mu(\mathbf{x})\mathop{}\!\mathrm{d}\pi^{\star}_{\varepsilon}(\mathbf{y}|\mathbf{x})\,.$
Knowing $\mu$, we can hence fully describe $\pi^{\star}_{\varepsilon}$ via the conditional distributions $(\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x}))_{\mathbf{x}\in\mathcal{X}}$. The latter are also of great practical interest, as they provide a way to transport a point $\mathbf{x}$ sampled from $\mu$ to the target domain $\mathcal{Y}$; either stochastically by sampling $\mathbf{y}_{1},...,\mathbf{y}_{n}$ from $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$, or deterministically by averaging over conditional samples: $T_{\varepsilon}(\mathbf{x}):=\mathbb{E}_{Y\sim\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})}[Y]$. Moreover, we can assess the uncertainty of these predictions using any statistic of $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$. This aspect is crucial for many real-world problems, especially in single-cell biology, as shown in section 5.1.
Figure 1: Sketch of the GENOT model. For each ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{x}}$ in the support of the source ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\bm{\mu}}$, we learn a flow $\phi_{1}(\cdot|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{x}})$ from the noise ${\color[rgb]{0.133,0.545,0.133}\definecolor[named]{pgfstrokecolor}{rgb}{0.133,0.545,0.133}\pgfsys@color@rgb@stroke{0.133}{0.545}{0.133}\pgfsys@color@rgb@fill{0.133}{0.545}{0.133}\bm{\rho}}$ to the conditional $\pi^{\star}_{\varepsilon}(\cdot|{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{x}})$, whose support lies in that of the target $\bm{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\pgfsys@color@rgb@stroke{0}{0}{1}\pgfsys@color@rgb@fill{0}{0}{1}\nu}}$. We highlight this procedure for a specific point ${\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\pgfsys@color@rgb@stroke{1}{0}{0}\pgfsys@color@rgb@fill{1}{0}{0}\mathbf{x}_{1}}$.
Learning the Conditional Distributions. Let $\rho=\mathcal{N}(0,I_{q})$ the standard Gaussian on the target space $\mathbb{R}^{q}\supset\mathcal{Y}$. From the noise outsourcing lemma (Kallenberg, 2002), there exists a collection of conditional generators $\{T^{\star}(\cdot|\mathbf{x})\}_{\mathbf{x}\in\mathcal{X}}$ s.t. $T^{\star}(\cdot|\mathbf{x}):\mathbb{R}^{q}\rightarrow\mathbb{R}^{q}$ and  for each $\mathbf{x}$ in the support of $\mu$, $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})=T^{\star}(\cdot|\mathbf{x})\sharp\rho$. This means that if $Z\sim\rho$, then $Y=T^{\star}(Z|\mathbf{x})\sim\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$. Here, we seek to learn a neural collection $\{T_{\theta}(\cdot|\mathbf{x})\}_{\mathbf{x}\in\mathcal{X}}$ of such conditional generators, fitting the constraint $T_{\theta}(\cdot|\mathbf{x})\sharp\rho=\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$, for any $\mathbf{x}$ in the support of $\mu$. We employ the CFM framework (see § 2) and parameterize each $T_{\theta}(\cdot|\mathbf{x})$ implicitly, as the flow induced by a neural vector field $v_{t,\theta}(\cdot|\mathbf{x}):\mathbb{R}^{q}\rightarrow\mathbb{R}^{q}$. Namely, $T_{\theta}(\cdot|\mathbf{x})=\phi_{1}(\cdot|\mathbf{x}):\mathbb{R}^{q}\to\mathbb{R}^{q}$ where $\phi_{t}(\cdot|\mathbf{x})$ solves
$\frac{\mathop{}\!\mathrm{d}}{\mathop{}\!\mathrm{d}t}\phi_{t}(\mathbf{z}|\mathbf{x})=v_{t,\theta}(\phi_{t}(\mathbf{z}|\mathbf{x})|\mathbf{x}),\quad\phi_{0}(\mathbf{z}|\mathbf{x})=\mathbf{z}.$ (3)
We stress that while $\mathbf{x}\in\mathcal{X}\subset\mathbb{R}^{d}$, the flow $\phi_{1}(\cdot|\mathbf{x})$ from $\rho$ to $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$ is defined on the target space $\mathbb{R}^{q}\supset\mathcal{Y}$. Hence, we can map samples within the same space when $p=q$, but also across incomparable spaces when $p\neq q$. In particular, this allows us to consider the quadratic OT problem (QEOT). Thus, for each $\mathbf{x}$, we optimize $v_{t,\theta}(\cdot|\mathbf{x})$ by minimizing the CFM loss (2) to map $\rho$ to $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$, i.e.
$\begin{split}\mathbb{E}_{t,Z\sim\rho,Y\sim\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})}[\|v_{t,\theta}\left([Z,Y]_{t}|\mathbf{x}\right)-(Y-Z)\|_{2}^{2}]\,.\end{split}$ (4)
where $[Z,Y]_{t}=(1-t)Z+tY$ interpolates between noise and conditional vectors. Averaging for all $\mathbf{x}$ in the support of $\mu$ and using Fubini’s Theorem, we derive $\mathcal{L}_{\mathrm{GENOT}}(\theta)$:
$\begin{split}\mathbb{E}_{t,Z\sim\rho,(X,Y)\sim\pi^{\star}_{\varepsilon}}[\|v_{t,\theta}\left([Y,Z]_{t}|X\right)-(Y-Z)\|_{2}^{2}]\,.\end{split}$ (5)
In practice, we optimize a sample-based GENOT loss by estimating $\pi^{n}_{\varepsilon}=\sum_{i,j}\mathbf{P}_{\varepsilon}^{ij}\delta_{(X_{i},Y_{j})}$, from $X_{1},\dots,X_{n}\sim\mu$ and $Y_{1},\dots,Y_{n}\sim\nu$, with a discrete EOT solver, see Alg. 1.
GENOT Preserves the Coupling Information, in the sense that, in the infinite sample regime, one could provably recover the original entropic coupling, whether in the linear or quadratic OT settings, and for any cost function.
Proposition 3.1 (Well-posedness of GENOT.).
Suppose that $\mathcal{L}_{\mathrm{GENOT}}(\theta)=0$. Then, for $X\sim\mu$, $Z\sim\rho$ and $Y=\phi_{1}(Z|X)$ denoting the solution of ODE (3), one has $Y\sim\pi^{\star}_{\varepsilon}(\cdot|X)$, and hence $(X,Y)\sim\pi^{\star}_{\varepsilon}$.
This result is noteworthy since standard flow matching only preserves straight couplings (Liu et al., 2022), while bridge matching (its stochastic counterpart) only preserves the linear EOT coupling for $c(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$, as shown recently by Bortoli et al. (2023). In contrast, GENOT is a conditional CFM model, learning the conditional distributions of the coupling independently of each other: For each $\mathbf{x}$, we leverage CFM to learn a conditional flow $\phi_{1}(\cdot|\mathbf{x})$ that maps $\rho$ to $\pi_{\varepsilon}^{\star}(\cdot|\mathbf{x})$. In this sense, GENOT can be conceptualized as an augmented flow matching procedure. We visualize that GENOT preserves the coupling information in Fig. 2.
GENOT Handles Any Cost. As GENOT preserves the coupling information, we can use it to approximate linear or quadratic EOT couplings. In both cases, we make no assumptions on the cost functions: We only need to evaluate them on samples to estimate $\pi^{\star}_{\varepsilon}$ using a discrete solver, see GENOT Alg. 1, line 5. In particular, we can use implicitly defined costs, whose evaluation requires a non-differentiable sub-routine. For instance, we can use the geodesic distance on the data manifold, which can be approximated from the shortest path distance on the $k$-nn graph induced by the Euclidean distance (Crane et al., 2013; Solomon et al., 2015). Using such a data-driven cost function in discrete OT pipelines has demonstrated its effectiveness in a range of single-cell genomics tasks (Demetci et al., 2022; Huguet et al., 2022; Klein et al., 2023), as comparing cells using $\ell_{p}$ distances often fails to capture discrepancies in gene expression space (Moon et al., 2018). With GENOT, we extend the use of this cost to the neural setting, for both linear (see § 5.1) and quadratic OT (see § 5.2).
GENOT Approximates Conditional Densities. For each $\mathbf{x}$, we build a CNF $p_{t}(\cdot|\mathbf{x})=\phi_{t}(\cdot|\mathbf{x})\sharp\rho$ between $p_{0}=\rho$ and $p_{1}=\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$. From the instantaneous change of variables formula (Chen et al., 2018), we can then approximate the conditional density $\pi^{\star}_{\varepsilon}(\mathbf{y}|\mathbf{x})$ at an arbitrary point $\mathbf{y}\in\mathcal{Y}$. We use it, for instance, in Figures 3, 6, LABEL:fig:gw_toy_data_densities and LABEL:fig:FGW_modality_translation_cond_density.
Figure 2: Fitting the EOT coupling between two Gaussian mixtures for the Coulomb cost (Benamou et al., 2015) $c(\mathbf{x},\mathbf{y})=1/\|\mathbf{x}-\mathbf{y}\|_{2}$, and $\varepsilon=0.01$, using GENOT and OT-CFM (Tong et al., 2023b). We connect paired samples with a line. The EOT coupling pairs source samples and target samples diagonally (left). GENOT (middle) generates samples correctly, while OT-CFM (right) fails to preserve the information from mini-batch couplings. The data-setting is inspired by Bortoli et al. (2023, Fig. 1)
3.2 U-GENOT: Extension to the Unbalanced Setting
Re-Balancing the UOT Problems. In its standard form, GENOT respects marginal constraints, so it cannot directly handle unbalanced formulations (ULEOT) or (UQEOT). We show that unbalanced EOT problems can be re-balanced. Eyring et al. (2023); Lübeck et al. (2022); Yang & Uhler (2019) introduced previously these ideas in the Monge map estimation setting, namely, in a static and deterministic setup. Our method stems from the fact that, for both linear and quadratic OT, the unbalanced EOT coupling $\pi^{\star}_{\varepsilon,\tau}$ between $\mu\in\mathcal{M}^{+}(\mathcal{X})$ and $\nu\in\mathcal{M}^{+}(\mathcal{Y})$ solves a balanced EOT problem between its marginals, which are re-weighted versions $\tilde{\mu}$ and $\tilde{\nu}$ of $\mu$ and $\nu$, that have the same mass.
Proposition 3.2 (Re-Balancing the unbalanced problems.).
Let $\pi^{\star}_{\varepsilon,\tau}$ be an unbalanced EOT coupling, solution of (ULEOT) or (UQEOT) between $\mu\in\mathcal{M}^{+}(\mathcal{X})$ and $\nu\in\mathcal{M}^{+}(\mathcal{Y})$. We note $\tilde{\mu}=p_{1}\sharp\pi^{\star}_{\varepsilon,\tau}$ and $\tilde{\nu}=p_{2}\sharp\pi^{\star}_{\varepsilon,\tau}$ its marginals. Then, in both cases, $\tilde{\mu}$ (resp. $\tilde{\nu}$) has a density w.r.t $\mu$ (resp. $\nu$) i.e. it exists $\eta,\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}$ s.t. $\tilde{\mu}=\eta\cdot\mu$ and $\tilde{\nu}=\xi\cdot\nu$. Moreover, $\tilde{\mu}$ and $\tilde{\nu}$ have the same mass and
  1. 1.
    (Linear) $\pi^{\star}_{\varepsilon,\tau}$ solves the balanced problem (LEOT) between $\tilde{\mu}$ and $\tilde{\nu}$ with the same $\varepsilon$.
  2. 2.
    (Quadratic) Provided that $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ (or $-c_{\mathcal{X}}$ and $-c_{\mathcal{Y}}$) are conditionally positive kernels (see Def. A.1), $\pi^{\star}_{\varepsilon,\tau}$ solves the balanced problem (QEOT) between $\tilde{\mu}$ and $\tilde{\nu}$ with $\varepsilon^{\prime}=m(\pi^{\star}_{\varepsilon,\tau})\,\varepsilon$, where $m(\pi^{\star}_{\varepsilon,\tau})=\pi^{\star}_{\varepsilon,\tau}(\mathcal{X}\times\mathcal{Y})$.
Learning the Coupling and the Re-Weightings Simultaneously. Thanks to Prop. 3.2, we aim to [i] learn a balanced EOT coupling between $\tilde{\mu}$ and $\tilde{\nu}$ along with [ii] the re-weighting functions $\eta,\xi$. Learning them is desirable since they model the creation and destruction of mass. We do both simultaneously by adapting the GENOT procedure. More formally, we seek to optimize $\mathcal{L}_{\mathrm{U\textrm{-}GENOT}}(\theta)$:
$\begin{split}&\mathbb{E}_{t,Z\sim\rho,(X,Y)\sim\pi^{\star}_{\varepsilon,\tau}}[\|v_{t,\theta}\left([Z,Y]_{t}|X\right)-(Y-Z)\|_{2}^{2}]\,\,\textrm{[i]}\\ &+\mathbb{E}_{X\sim\mu}[(\eta-\eta_{\theta})(X)^{2}]+\mathbb{E}_{Y\sim\nu}[(\xi-\xi_{\theta})(Y)^{2}]\,\,\,\,\,\,\,\textrm{[ii]}\end{split}$
As with GENOT, we simply need to estimate a discrete unbalanced EOT coupling $\pi_{\varepsilon,\tau}^{n}$ from samples to estimate that loss. We build upon theoretical insights on the linear OT case and extend them to the quadratic OT case in practice.
Proposition 3.3 (Poitwise estimation of re-weighting functions.).
Let $\pi_{\varepsilon,\tau}^{n}=\sum_{i,j}\mathbf{P}_{\varepsilon,\tau}^{i,j}\delta_{(X_{i},Y_{j})}$, solution to (ULEOT) between empirical counterparts of $\mu$ and $\nu$. Let $\mathbf{a}=\mathbf{P}_{\varepsilon,\tau}\mathbf{1}_{n}$ and $\mathbf{b}=\mathbf{P}_{\varepsilon,\tau}^{\top}\mathbf{1}_{n}$ its marginals weights. Then, almost surely, $n\,a_{i}\to\eta(X_{i})$ and $n\,b_{i}\to\xi(Y_{i})$.
Using Prop. 3.2, $\hat{\pi}_{\varepsilon,\tau}^{n}$ is a balanced EOT coupling between its marginals, which are empirical approximations of $\tilde{\mu}$ and $\tilde{\nu}$. Hence, we estimate the term [i] of the loss as we do in the balanced case by sampling from the discrete coupling. Furthermore, Prop.3.3 highlights that the estimation of $\hat{\pi}_{\varepsilon,\tau}^{n}$ also provides a consistent estimate of the re-weighting function evaluations at each $X_{i}$ and $Y_{i}$. This enables estimating [ii]. Therefore, switching from GENOT to U-GENOT simply involves using an unbalanced solver instead of a balanced one, and regressing the re-weightings on the marginal weights of the estimated discrete coupling. We detail our procedure in Alg. 1, showing the additional steps w.r.t. GENOT in teal.
3.3 Combining Linear and Quadratic OT
We show in § 3.1 and § 3.2 how to use GENOT to solve OT problems within the same space or across incomparable spaces. On the other hand, numerous real-world problems pose the challenge of the source and target domains being only partially incomparable (Vayer et al., 2018). Therefore, suppose that the source and target space can be decomposed as $\mathcal{X}=\Omega\times\mathcal{\tilde{X}}$ and $\mathcal{Y}=\Omega\times\mathcal{\tilde{Y}}$, respectively. Intuitively, a sample $(\mathbf{u},\mathbf{x})\in\Omega\times\mathcal{\tilde{X}}$ can be interpreted as a structural information $\mathbf{x}$ equipped with a feature $\mathbf{u}$. Afterwards, assume we are given a cost $c:\Omega\times\Omega\rightarrow\mathbb{R}$ to compare features, along with the intra-domain costs $c_{\mathcal{\tilde{X}}},c_{\mathcal{\tilde{Y}}}$. The entropic Fused Gromov-Wasserstein (FGW) problem reads
$\min_{\pi\in\Pi(\mu,\nu)}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}Q_{c_{\tilde{\mathcal{X}}},c_{\tilde{\mathcal{Y}}}}^{c}\mathop{}\!\mathrm{d}(\pi\otimes\pi)+\varepsilon\mathrm{KL}(\pi|\mu\otimes\nu)\,,$ EFGW
where $Q_{c_{\tilde{\mathcal{X}}},c_{\tilde{\mathcal{Y}}}}^{c}\left((\mathbf{u},\mathbf{x}),(\mathbf{v},\mathbf{y}),\mathbf{x}^{\prime},\mathbf{y}^{\prime}\right):=(1-\alpha)\,c(\mathbf{u},\mathbf{v})\,+\,\alpha\,|c_{\mathcal{\tilde{X}}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{\tilde{Y}}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}$ and $\alpha\in[0,1]$. This loss combines the pointwise structural distortion and the feature information. The fused problem (EFGW) also admits an unbalanced extension, which can be derived in the same way as (UQEOT) using the quadratic $\mathrm{KL}^{\otimes}$ (Thual et al., 2023).
(U-)GENOT for the Fused Setting. Whether in the balanced or unbalanced setting, we can use our method to learn a specific coupling as soon as it can be estimated from samples. We stress that the discrete solvers we use for (QEOT) and (UQEOT), respectively Peyré et al. (2016)’s and Séjourné et al. (2023b)’s ones, are still applicable in the fused setting. As a result, we can approximate solutions of (EFGW) and its unbalanced counterpart with (U-) GENOT. To illustrate the learning outcome, take a solution $\pi^{\star}_{\alpha}$ of (EFGW). Learning $\pi^{\star}_{\alpha}$ with our method amounts to training vector fields that are conditioned on pairs of modalities from the source domain $v_{t,\theta}(\cdot,|\mathbf{u},\mathbf{x})$, to sample pairs of modalities from the target domain via the flow: $Z\sim\rho$, $\phi_{1}(Z|\mathbf{u},\mathbf{x})=(V,Y)\sim\pi^{\star}_{\alpha}(\cdot|\mathbf{u},\mathbf{x})$. The sampled modalities $(V,Y)$ (i) minimize transport cost quantified by $c$ along the feature (ii) while minimizing the distortion along the structural information quantified by $c_{\tilde{\mathcal{X}}}$ and $c_{\tilde{\mathcal{Y}}}$. In § 5.2, we use fused couplings to enhance the translation between different modalities of cells by adding label information .
Algorithm 1 U-GENOT. Skip teal steps for GENOT.
1:  Require parameters: Batch size $n$; entropic regularization $\varepsilon$; unbalancedness parameter $\tau$; linear, quadratic or fused discrete OT solver $\mathrm{Solver}_{\varepsilon,{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\pgfsys@color@rgb@stroke{0}{.5}{.5}\pgfsys@color@rgb@fill{0}{.5}{.5}\tau}}$.
2:  Require network: Time-dependent conditional velocity field $v_{t,\theta}(\cdot|\cdot):\mathbb{R}^{q}\times\mathbb{R}^{p}\to\mathbb{R}^{q}$; re-weighting functions $\eta_{\theta}:\mathbb{R}^{p}\to\mathbb{R}$, $\xi_{\theta}:\mathbb{R}^{q}\to\mathbb{R}$.
3:  for $t=1,\dots,T_{\mathrm{iter}}$ do
4:     Sample $X_{1},\dots,X_{n}\sim\mu$ and $Y_{1},\dots,Y_{n}\sim\nu$.
5:      $\mathbf{P}_{\varepsilon,{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\pgfsys@color@rgb@stroke{0}{.5}{.5}\pgfsys@color@rgb@fill{0}{.5}{.5}\tau}}\leftarrow\mathrm{Solver}_{\varepsilon,{\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\pgfsys@color@rgb@stroke{0}{.5}{.5}\pgfsys@color@rgb@fill{0}{.5}{.5}\tau}}\left(\{X_{i}\}_{i=1}^{n},\{Y_{i}\}_{i=1}^{n}\right)\in\mathbb{R}_{+}^{n\times n}$.
6:     $\mathbf{a}\leftarrow\mathbf{P}_{\varepsilon,\tau}\mathrm{1}_{n}$ and $\mathbf{b}\leftarrow\mathbf{P}_{\varepsilon,\tau}^{\top}\mathrm{1}_{n}$.
7:      Sample $(i_{1},j_{1}),...,(i_{n},j_{n})\sim\mathbf{P}_{\varepsilon,\tau}$.
8:     Sample $Z_{1},...,Z_{n}\sim\rho$ and $t_{1},...,t_{n}\sim\mathcal{U}([0,1])$.
9:     $\mathcal{L}(\theta)\leftarrow\sum_{k}\|v_{t,\theta}([Z_{k},Y_{j_{k}}]_{t}|X_{i_{k}})-(Y_{j_{k}}-Z_{k})\|_{2}^{2}$
10:       ${\color[rgb]{0,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{0,.5,.5}\pgfsys@color@rgb@stroke{0}{.5}{.5}\pgfsys@color@rgb@fill{0}{.5}{.5}+\sum_{k}(\eta_{\theta}(X_{k})-n\mathbf{a}_{k})^{2}+(\xi_{\theta}(Y_{k})-n\mathbf{b}_{k})^{2}}$.
11:     $\theta\leftarrow\textrm{Update}(\theta,\tfrac{1}{n}\nabla\mathcal{L}(\theta))$.
12:  end for
4 Related work and Discussions
Static Neural EOT. While GENOT is the first model to consider the quadratic (and Fused) EOT setting, as well as addressing unbalancedness, various methods have been proposed in the (balanced) linear EOT scenario. The first class of methods solves (LEOT) dual. While some of them (Seguy et al., 2017; Genevay et al., 2019a) do not allow direct sampling from to $\pi^{\star}_{\varepsilon}$aniels et al. (2021) and Mokrov et al. (2023) model the conditional distributions $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$. However, these methods might be costly and unstable, as they rely on Langevin sampling during training or inference.
Dynamic Neural EOT. The second class of linear EOT solvers builds upon the link between (LEOT) and the SB problem  (De Bortoli et al., 2021; Chen et al., 2021; Vargas et al., 2021; Gushchin et al., 2022). As simulations are costly in this setting, recent works consider simulation-free training via bridge matching (Shi et al., 2023; Korotin et al., 2023a; Liu et al., 2023; Zhou et al., 2023; Peluchetti, 2023; Tong et al., 2023a, b). While Tong et al. (2023a, b), use mini-batch OT similarly to us, our method is fundamentally different, as we do not build upon the link between EOT and SB. We only use flow matching as a powerful black box to learn a flow from the noise $\rho$ to each conditional $\pi_{\varepsilon}^{\star}(\cdot|\mathbf{x})$. Therefore, our approach allows for more flexibility. First (i), the abovementioned methods assume that $\mathcal{X},\mathcal{Y}\subset\mathbb{R}^{d}$, since they learn a velocity field (or a drift) by directly bridging $\mathbf{x}\in\mathcal{X},\mathbf{y}\in\mathcal{Y}$. Then, they map $\mu$ to $\nu$ with the induced (stochastic) flow. On the other hand, conditionally to each $\mathbf{x}\in\mathcal{X}\subset\mathbb{R}^{d}$, we learn a velocity field $v_{t,\theta}(\cdot|\mathbf{x}):\mathbb{R}^{q}\to\mathbb{R}^{q}$ in the target space, by building paths between noise $\mathbf{z}\in\mathbb{R}^{q}$ and $\mathbf{y}\in\mathcal{Y}\subset\mathbb{R}^{q}$. Then, we map $\rho$ to each $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$ with the flows $\phi_{1}(\cdot|\mathbf{x})$, and recover $\nu=\phi_{1}(\cdot|\cdot)\sharp(\rho\otimes\mu)$. Second (ii), since they use discrete couplings to pair the samples $\mathbf{x},\mathbf{y}$ that they bridge, they can only approximate the SB for $c(\mathbf{x},\mathbf{y})=\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$. This results from (Bortoli et al., 2023, Prop. 2). In contrast, our method handles any cost, as shown in Fig 2. Third (iii), as they learn a single flow directly transporting $\mu$ to $\nu$, they do not approximate the conditional densities $\pi_{\varepsilon}^{\star}(\mathbf{y}|\mathbf{x})$. Similarly to Tong et al. (2023b), Pooladian et al. (2023) couple samples from $\mu$ and $\nu$, but they only model deterministic maps and assume $\mu=\mathcal{N}(0,I_{d})$. Finally, Bortoli et al. (2023) have recently proposed an augmented bridge matching procedure that preserves couplings. Yet, they still require $\mathcal{X},\mathcal{Y}\subset\mathbb{R}^{d}$ and do not approximate conditional densities.
Mini-batches and Biases Quantifying non-asymptotically the bias resulting from minimizing a sample-based GENOT loss, and not its population value, is challenging. The OT curse of dimensionality (Weed & Bach, 2017) has been discussed in generative models relying on mini-batch couplings (Genevay et al., 2019a; Fatras et al., 2021; Tong et al., 2023b, a). Yet, our goal is not to model non-regularized OT, such as a deterministic Monge map, or a Benamou-Brenier vector field. We explicitly target the entropy-regularized OT coupling. Thus, using $\varepsilon\gg 0$ helps to mitigate the curse of dimensionality because of two qualitative factors:
  • (i)
    Statistical. For both linear and quadratic OT, all statistical recovery rates that relate to entropic costs (Genevay et al., 2019b; Zhang et al., 2023) or maps (Pooladian & Niles-Weed, 2021; Rigollet & Stromme, 2022), have a far more favorable regime, with a parametric rate in $\varepsilon>0$ that dodges the curse of dimensionality.
  • (ii)
    Computational. While the benefits of employing a large enough $\varepsilon$ in Sinkhorn’s algorithm are widely known, Rioux et al. (2023) have recently shown that as $\varepsilon$ increases, the quadratic OT problem (QEOT) becomes convex, making discrete solvers faster and more reliable.
To demonstrate this aspect, we empirically study the influence of the batch size on GENOT, using data from a recent benchmark (Gushchin et al., 2023), see Fig. LABEL:fig:eot_benchmark_all.
Figure 3: Estimating the  (ULEOT) coupling between Gaussian mixtures with U-GENOT-L ($\varepsilon=0.05$, $\tau_{1}=\tau_{2}=0.98$ and $c=\ell_{2}^{2}$). Left to right: Source (top) $\mu$ and the target $\nu$ (bottom), learned (top) and true (bottom) coupling, learned (top) and true (bottom) re-weighted marginal, and conditional densities. We further showcase the influence of $\tau_{1},\tau_{2}$ in Fig. LABEL:fig:1d_simulated_unbalanced_big.
5 Experiments
Model rank $\operatorname{B\mathbb{W}_{2}^{2}-UVP}$ rank $\operatorname{cB\mathbb{W}_{2}^{2}-UVP}\downarrow$
LSOT 9.33 9.33
SCONES 7.50 7.67
DiffSB 7.58 7.50
FB-SDE-J 5.12 7.33
FB-SDE-A 6.75 6.75
EgNOT 2.71 5.75
ENOT 6.08 3.00
NOT 4.92 2.92
MLE-SB 3.17 2.67
GENOT-L 1.79 2.08
Table 1: Average ranks of each method across experiments, on estimating the ground-truth EOT coupling $\pi^{\star}_{\varepsilon}$ between each pair of measures $\mu,\nu$, in dimension $d\in\{2,16,64,128\}$ and $\varepsilon\in\{0.1,1,10\}$ of Gushchin et al. (2023)’s benchmark. As them, we measure performances with both $\operatorname{B\mathbb{W}_{2}^{2}-UVP}$ (LABEL:eq:bures_wasserstein) and $\operatorname{cB\mathbb{W}_{2}^{2}-UVP}$ (LABEL:eq:conditional_bures_wasserstein) metrics. See App. LABEL:app:eot_benchmark for details.
Figure 4: Top: Source cell from the early time points (top left) and samples of the conditional distributions of the EOT coupling learned with GENOT for the geodesic cost $d_{\mathcal{M}}$ (middle) and the $\ell_{2}^{2}$ cost (right) projected onto a UMAP (McInnes et al., 2018). Using the geodesic distance as cost function prevents $\pi_{\theta}$ from skipping the intermediate cell population, which is far away from the source cell and terminal cells in $\ell_{2}$ distance. Bottom: UMAP colored according to the uncertainty score $\mathrm{cos}\text{-}\mathrm{var}(\pi_{\theta}^{d_{\mathcal{M}}}(\cdot|X))$ of each source cell $\mathbf{x}$. Target cells are colored in gray.
We show the applicability and versatility of our framework to map within the same space (solving a linear EOT problem) and across incomparable spaces (solving a purely quadratic or fused EOT problem). Metrics and datasets are discussed in App. LABEL:app:metrics and App. LABEL:app:datasets, respectively. Further experimental details or results are reported in App. LABEL:app:experiments. Setups for baselines are listed in App. LABEL:app:competing_methods. Implementation details can be found in App. LABEL:app:implementation. We denote by GENOT-L the GENOT model for solving the linear problem (LEOT), GENOT-Q for the quadratic (QEOT) and GENOT-F for the fused (EFGW) one. When considering the unbalanced counterparts, we add the prefix U. Moreover, when using the conditional mean of a GENOT model, we add the suffix CM. We denote our estimated couplings by $\pi_{\theta}$.
5.1 GENOT for Linear EOT problems
GENOT-L on a static SB benchmark. First, we show that GENOT accurately computes EOT couplings using a recent benchmark. For different $\varepsilon\in\{0.1,1,10\}$ and $d\in\{2,16,64,128\}$), Gushchin et al. (2023) consider pairs $\mu,\nu$ for which the $\pi^{\star}_{\varepsilon}$ solution of (LEOT) with $\ell_{2}^{2}$ cost $c(\mathbf{x},\mathbf{y})=\frac{1}{2}\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$ is known. Tab. 1 shows that GENOT-L outperforms all baselines, choosing one configuration for all experiments as opposed to some baselines. Detailed results can be found in Tab. LABEL:table:bw_uvp and LABEL:table:cbw_uvp. Importantly, GENOT’s performance can be further improved when adapting the batch size, see Fig. LABEL:fig:eot_benchmark_all. These results are of particular significance since many baselines are tailored towards the specific task of solving (i) the linear EOT problem (LEOT), (ii) in the balanced setting (iii) with the $\ell_{2}^{2}$ cost. In contrast, the contributions of GENOT go beyond this restrictive setting. In the following, we demonstrate GENOT’s flexibility and highlight its relevance in various single-cell biology tasks.
U-GENOT-L on simulated data. To go beyond the balanced setting, Fig. 3 visualizes U-GENOT-L’s ability to learn unbalanced EOT plans and re-weighting functions. We further showcase the influence of $\tau_{1},\tau_{2}$ in Fig. LABEL:fig:1d_simulated_unbalanced_big.
GENOT-L recovers single cell trajectories. We demonstrate GENOT-L’s capability to infer the development of cells on a dataset capturing gene expression of the developing mouse pancreas at embryonic days 14.5 (source) and 15.5 (target) (Bastidas-Ponce et al., 2019). As cells are known to evolve stochastically (Elowitz et al., 2002), we leverage GENOT’s capability to generate samples from the conditional distribution to model non-deterministic trajectories. Moreover, as Euclidean metrics often fail to represent meaningful distances in gene expression space (Moon et al., 2018), we leverage GENOT’s flexibility and use the geodesic distance on the data manifold $d_{\mathcal{M}}$ as cost, as discussed in § 3.1. Here, $\mathcal{M}=\mathcal{X}\cup\mathcal{Y}$ since the source and target domains have the same dimension. We also estimate the EOT coupling for the $\ell_{2}^{2}$ cost for comparison. Fig. 4 and Fig. LABEL:fig:pancreas_transitions highlight that the geodesic cost causes the conditional distribution to follow the data manifold. In particular, this helps to obtain biologically more plausible cell trajectories. We follow Gayoso et al. (2022) for assessing the uncertainty of cell trajectories by computing $\mathrm{cos}\text{-}\mathrm{var}(\pi_{\theta}(\cdot|\mathbf{x}))=\operatorname{Var}_{Y\sim\pi_{\theta}(\cdot|\mathbf{x})}[\mathrm{cos}\text{-}\mathrm{sim}(Y,\mathbb{E}_{Y\sim\pi_{\theta}(\cdot|\mathbf{x})}[Y])]$ See App. LABEL:app:metrics_general for more details on $\mathrm{cos}\text{-}\mathrm{var}$. We expect high uncertainty in cell types in early developmental stages and low variance in mature cell types. Indeed, Fig. 4 and Fig. LABEL:fig:pancreas_uncertainty_suppl show that the computed uncertainty is biologically meaningful: GENOT-L helps to uncover lineage branching events.
Figure 5: Accuracy of cellular response predictions of U-GENOT-L for different cancer drugs with varying unbalancedness parameter $\tau=\tau_{1}=\tau_{2}$. As $\tau$ decreases, we allow for more unbalancedness. For each $\tau$, we display results on 3 runs.
U-GENOT-L predicts single-cell responses to perturbations. Neural OT has been successfully applied to model cellular responses to perturbations using deterministic maps (Bunne et al., 2021). GENOT has the comparative advantage to model conditional distributions, allowing for uncertainty quantification. We consider single cell RNA-seq data measuring the response of cells to 163 cancer drugs (Srivatsan et al., 2020). Each drug has been applied to a population of cells that can be partitioned into 3 different cell types. Source (resp. target) is cells before (resp. after) drug application. While there is no ground truth in the matching between unperturbed and perturbed cells due to the destructive nature of sequencing technologies, we know which unperturbed subset of cells is supposed to be mapped to which perturbed subset of cells. We use this to define an accuracy metric (see App. LABEL:app:metrics_cell_type_transition), while we assess the uncertainty of the prediction using $\mathrm{cos}\text{-}\mathrm{var}$. Fig. LABEL:fig:perturbation_calibration shows that for 117 out of 163 drugs, the model is perfectly calibrated (see App. LABEL:app:calibration_score), while it yields a negative correlation between error and uncertainty only for one drug. We can improve predictions by accounting for class imbalances between different cell types using U-GENOT. Indeed, Fig. 5 shows that allowing for mass variation increases accuracy for nine different cancer drugs, which we selected as they are known to have a strong effect. Fig. LABEL:fig:dacinostat and LABEL:fig:tanespimycin confirm the results visually.
Figure 6: Mapping a Swiss roll in $\mathbb{R}^{3}$ ($\mu$, top left) to a spiral in $\mathbb{R}^{2}$ ($\nu$, bottom left) with GENOT-Q. Center: Color code tracks where samples from $\mu$ (top) are mapped to (bottom). Right column: samples from $\mu$ (top) and each of the corresponding conditionals, along with conditional density estimates. The learned QEOT coupling minimizes the cost distortion: Points close in support of $\mu$ are paired to points close in the support of $\nu$.
Figure 7: Top: UMAP of transported cells with GENOT-F (colored by cell type) and cells in the target distribution (gray). GENOT-F clusters similar cells together, i.e., according to their cell type. Bottom: UMAP of transported cells with a GENOT model trained on batch-wise independent couplings, thus not using OT. This remains to replace $\pi^{\star}_{\varepsilon}$ by $\mu\otimes\nu$ in Eq. (5). While the learned coupling then randomly mixes the cell types, it also degrades the global fitting of the target distribution of cells.
Figure 8: Benchmark of GENOT-Q models against discrete GW (GW-LR, App. LABEL:app:competing_methods) on translating cells between ATAC space of dimension $d_{1}$ and RNA space of dimension $d_{2}$. Performance is measured with the FOSCTTM score (App. LABEL:app:metrics_single_cell) and the Sinkhorn divergence between target and predicted target distribution. (left) intra-domain costs $c_{\mathcal{X}}=c_{\mathcal{Y}}=\ell_{2}^{2}$, (right) geodesic distances $c_{\mathcal{X}}=d_{\mathcal{X}}$ and $c_{\mathcal{Y}}=d_{\mathcal{Y}}$. Mean and 95% CI shown across 3 runs.
5.2 GENOT for Quadratic EOT Problems
GENOT-Q on simulated data. We transport a Swiss role in $\mathbb{R}^{3}$ to a spiral in $\mathbb{R}^{2}$. Fig. 6 shows that GENOT-Q successfully mimics an isometric alignment. Here, we set $\varepsilon=0.01$ and investigate its influence in more detail in Fig. LABEL:fig:gw_toy_data_eps.
GENOT-Q translates modalities of single cells. The number of modalities of a cell that can be simultaneously measured is limited due to technical limitations. Yet, it is important to match measurements of different modalities to obtain a more holistic view of the profile of a cell. Demetci et al. (2022) address this task using discrete quadratic OT. However, due to the lack of neural quadratic EOT solvers, no OT-inspired method naturally allows for out-of-sample prediction. In the following, we use GENOT-Q to translate ATAC measurements (source) to gene expression space (target) on a bone marrow dataset (Luecken et al., 2021). As both modalities were measured in the same cell, the true match of each cell is known for this specific dataset. We compare GENOT-Q with discrete OT extended out-of-sample with linear regression (GW-LR, see LABEL:app:regression_discrete). We assess the performance using (i) the FOSCTTM (“Fractions of Samples Closer to the True Match”, see LABEL:app:foscttm) that measures the optimality of the learned coupling, and (ii) the Sinkhorn divergence (Feydy et al., 2019) between the predicted target and the target to assess distributional fitting. As in § 5.1, we leverage GENOT’s flexibility and use as intra-domains costs the geodesic distances on the source and target domain, namely $c_{\mathcal{X}}=d_{\mathcal{X}}$, $c_{\mathcal{Y}}=d_{\mathcal{Y}}$. We also estimate the EOT coupling for the $\ell_{2}^{2}$ cost for comparison. Results are shown in Fig. 8. Regarding the FOSCTTM score, we can see that (i) using geodesic costs is crucial in high dimensions and (ii) GW-LR is competitive in low dimensions but not for higher ones. Regarding distributional fitting, GENOT models outperform by orders of magnitude.
GENOT-F improves modality translation. To enhance the performances attained by purely quadratic OT-based models, we introduce a novel method for translating between ATAC and RNA. We extend the idea proposed by Demetci et al. (2022) to the fused setting: We approximate RNA data from ATAC measurements using gene activity (Stuart et al., 2021), and we further process the data using a conditional VAE (Lopez et al., 2018a) to reduce batch effects. This way, we construct a joint space $\Omega$. Following the notations in § 3.3, RNA (source $\mathbf{x}\in\tilde{\mathcal{X}}$) and ATAC (target $\mathbf{y}\in\tilde{\mathcal{Y}}$) carry the structural information, while features $\mathbf{u}\in\Omega$ and $\mathbf{v}\in\Omega$ are obtained from the VAE embedding. Fig. LABEL:fig:fgw_modality_translation_diff_alphas shows that the additional linear penalty on the feature term helps to obtain a better alignment compared to GENOT-Q. Fig. 7 visualizes the learned fused coupling compared to GENOT trained with batch-wise independent couplings. When aligning multiple modalities of single cells, the proportions of cell types in the two modalities might differ. We simulate this setting by removing cells belonging to certain cell types Tab. LABEL:table:unbalancedness_ugenot_fgw shows that U-GENOT-F preserves high accuracy while learning meaningful rescaling functions.
Conclusion. We introduce GENOT, a versatile neural OT framework to learn cost-efficient stochastic maps, within the same space, or across incomparable spaces. We can loosen the mass conservation constraint if desired, while transport plans can be learned using any cost. We demonstrate that GENOT’s flexibility meets numerous requirements necessary for various prominent tasks in single-cell biology. GENOT models are implemented in ott-jax (Cuturi et al., 2022).
Impact Statement. This work introduces novel approaches in neural optimal transport and presents applications in single-cell biology. Leveraging single-cell genomics data holds promise for advancing personalized medicine but requires cautious handling due to its potential inclusion of sensitive information. Moreover, as neural optimal transport techniques find applicability across diverse domains, the societal ramifications of this research extend beyond single-cell genomics. With our publicly available implementation of GENOT in ott-jax (Cuturi et al., 2022), we enable its utilization by general machine learning practitioners. In a planned future release in moscot (Klein et al., 2023), we ensure GENOT’s accessibility to advance and accelerate research specifically tailored for single-cell biologists.
Acknowledgement. Co-funded by the European Union (ERC, DeepCell - 101054957). Views and opinions expressed are, however, those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. F.J.T. consults for Immunai Inc., Singularity Bio B.V., CytoReason Ltd, Cellarity, and has ownership interest in Dermagnostix GmbH and Cellarity.
Appendix
Appendix A Proofs
A.1 Proofs of § 1
See 3.1
Proof.
This result follows directly from the construction of the loss. Suppose that $\mathcal{L}_{\mathrm{GENOT}}(\theta)=0$. By Fubini’s Theorem, this implies that:
$\begin{split}&\mathbb{E}_{X\sim\mu}\mathbb{E}_{t\sim\mathcal{U}([0,1]),Z\sim\rho,Y\sim\pi^{\star}_{\varepsilon}(\cdot|X)}[\|v_{t,\theta}\left([Z,Y]_{t}|X\right)-(Y-Z)\|_{2}^{2}]=0\,.\\ \textrm{i.e.}\quad&\mathbb{E}_{X\sim\mu}\mathbb{E}_{t\sim\mathcal{U}([0,1]),Z\sim\rho,Y\sim\pi^{\star}_{\varepsilon}(\cdot|X)}[\|v_{t,\theta}\left(tZ+(1-t)Y|X\right)-(Y-Z)\|_{2}^{2}]=0\,.\end{split}$ (6)
Since it is positive, the function
$\ell:\mathbf{x}\mapsto\mathbb{E}_{t\sim\mathcal{U}([0,1]),Z\sim\rho,Y\sim\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})}[\|v_{t,\theta}\left(tZ+(1-t)Y|\mathbf{x}\right)-(Y-Z)\|_{2}^{2}]$
is therefore zero $\mu$-a.e., which means that for any $\mathbf{x}$ in the support of $\mu$, one has $\ell(\mathbf{x})=0$. Moreover, $\ell(\mathbf{x})$ is the CFM loss with source $\rho$ and target $\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$, applied to the vector field $v_{t,\theta}(\cdot|\mathbf{x})$. Therefore, $\ell(\mathbf{x})=0$ implies that $v_{t,\theta}(\cdot|\mathbf{x})$ generates a probability path between $\rho$ and $\pi^{\star}(\cdot|\mathbf{x})$ thanks to (Lipman et al., 2023, Theorem 1). This means that if $\phi_{t}(\cdot|\mathbf{x})$ is the flow induced by $v_{t,\theta}(\cdot|\mathbf{x})$, one has $\phi_{1}(\cdot|\mathbf{x})\sharp\rho=\pi^{\star}_{\varepsilon}(\cdot|\mathbf{x})$. Afterwards, if $X\sim\mu$, $Z\sim\rho$ and $Y=\phi_{1}(Z|X)$ denotes the solution of ODE (3), one has $Y\sim\pi^{\star}_{\varepsilon}(\cdot|X)$, so $(X,Y)\sim\pi^{\star}_{\varepsilon}$ by definition. ∎
See 3.2 We first remind the definition of a conditionally positive kernel, which is involved in point 2. of Prop. 3.2.
Definition A.1.
A kernel $k:\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ is conditionally positive if it is symmetric and for any $\mathbf{x}_{1},...,\mathbf{x}_{n}\in\mathbb{R}^{d}$ and $\mathbf{a}\in\mathbb{R}^{n}$ s.t. $\mathbf{a}^{\top}\mathbf{1}_{n}=0$, one has
$\sum_{i,j=1}^{n}a_{i}a_{j}\,k(\mathbf{x}_{i},\mathbf{x}_{j})\geq 0$
Conditionally positive kernels include all positive kernels, such $k(\mathbf{x},\mathbf{y})=\langle\mathbf{x},\mathbf{y}\rangle$ or $k_{\gamma}(\mathbf{x},\mathbf{y})=\exp(-\tfrac{1}{\gamma}\|\mathbf{x}-\mathbf{y}\|_{2}^{2})$, but also the negative squared Euclidean distance $k(\mathbf{x},\mathbf{y})=-\|\mathbf{x}-\mathbf{y}\|_{2}^{2}$.
Proof of 3.2.
Step 1: Re-weightings. We first show that for $\pi^{\star}_{\varepsilon,\tau}$ solution of (LEOT) or (QEOT), it exists $\eta,\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{+}$ s.t. $\tilde{\mu}=p_{1}\sharp\pi^{\star}_{\varepsilon,\tau}=\eta\cdot\mu$ and $\tilde{\nu}=p_{1}\sharp\pi^{\star}_{\varepsilon,\tau}=\xi\cdot\nu$.
We start with the linear case and consider $\pi^{\star}_{\varepsilon,\tau}$ solution of (LEOT). The result follows from duality. Indeed, from (Séjourné et al., 2023a, Prop. 2), one has the existence of the so-called entropic potentials $f^{\star}\in\mathcal{C}(\mathcal{X})$ and $g^{\star}\in C(\mathcal{Y})$ s.t.
$\frac{\mathop{}\!\mathrm{d}\pi^{\star}_{\varepsilon,\tau}}{\mathop{}\!\mathrm{d}(\mu\otimes\nu)}(\mathbf{x},\mathbf{y})=\exp\left(\frac{f^{\star}(\mathbf{x})+g^{\star}(\mathbf{y})-c(\mathbf{x},\mathbf{y})}{\varepsilon}\right)$ (7)
Therefore, $\tilde{\mu}=\eta\cdot\mu$ and $\tilde{\nu}=\xi\cdot\nu$ where $\eta:\mathcal{X}\to\mathbb{R}^{+}$ and $\xi:\mathcal{Y}\to\mathbb{R}^{+}$ are defined by:
$\begin{split}&\eta(\mathbf{x})=\int_{\mathcal{Y}}\exp\left(\frac{f^{\star}(\mathbf{x})+g^{\star}(\mathbf{y})-c(\mathbf{x},\mathbf{y})}{\varepsilon}\right)\mathop{}\!\mathrm{d}\nu(\mathbf{y})\quad\textrm{and}\quad\xi(\mathbf{y})=\int_{\mathcal{X}}\exp\left(\frac{f^{\star}(\mathbf{x})+g^{\star}(\mathbf{y})-c(\mathbf{x},\mathbf{y})}{\varepsilon}\right)\mathop{}\!\mathrm{d}\mu(\mathbf{x})\,.\end{split}$ (8)
We now handle the quadratic case and consider $\pi^{\star}_{\varepsilon,\tau}$ solution of (UQEOT). We remind that the (UQEOT) problem between $\mu$ and $\nu$ reads:
$\begin{split}\min_{\pi\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\,+\varepsilon\mathrm{KL}^{\otimes}(\pi|\mu\otimes\nu)+\lambda_{1}\mathrm{KL}^{\otimes}(\pi_{1}|\mu)+\lambda_{2}\mathrm{KL}^{\otimes}(\pi_{2}|\nu),\end{split}$
For $\pi\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})$, the quadratic relative entropy $\mathrm{KL}^{\otimes}(\pi|\mu\otimes\mu)=\mathrm{KL}(\pi\otimes\pi|(\mu\otimes\nu)\otimes(\mu\otimes\nu))$ is finite i.f.f. $\pi\otimes\pi$ has a density w.r.t. $(\mu\otimes\nu)\otimes(\mu\otimes\nu)$, which implies that $\pi$ has a density w.r.t. $\mu\otimes\nu$. Therefore, one can reformulate (UQEOT) as:
$\begin{split}\min_{h\in L_{1}^{+}(\mathcal{X}\times\mathcal{Y})}\int_{(\mathcal{X}\times\mathcal{Y})^{2}}&\left(c_{\mathcal{X}}(\mathbf{x},\mathbf{x})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y})\right)^{2}h(\mathbf{x},\mathbf{y})h(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\mathop{}\!\mathrm{d}\mu(\mathbf{x})\mathop{}\!\mathrm{d}\mu(\mathbf{x}^{\prime})\mathop{}\!\mathrm{d}\nu(\mathbf{y})\mathop{}\!\mathrm{d}\nu(\mathbf{y}^{\prime})\\ &+\varepsilon\mathrm{KL}(h|\mu\otimes\nu)+\tau_{1}\mathrm{KL}(h_{1}|\mu)+\tau_{2}\mathrm{KL}(h_{2}|\nu)\end{split}$ (9)
where we extend the KL divergence for densities: $\mathrm{KL}(r|\gamma)=\int(r\log(r)+r-1)\mathop{}\!\mathrm{d}\gamma$, and define the marginal densities $h_{1}:\mathbf{x}\in\mathcal{X}\mapsto\int_{\mathcal{Y}}h(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\nu(\mathbf{y})$ and $h_{2}:\mathbf{y}\in\mathcal{Y}\mapsto\int_{\mathcal{X}}h(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\nu(\mathbf{y})$. As a result, it exists $h^{\star}\in L_{1}^{+}(\mathcal{X}\times\mathcal{Y})$ s.t. $\pi^{\star}_{\varepsilon,\tau}=h^{\star}\cdot\mu\otimes\nu$. It follows that $\tilde{\mu}=\eta\cdot\mu$ and $\tilde{\nu}=\xi\cdot\nu$ with $\eta=h^{\star}_{1}$ and $\xi=h^{\star}_{2}$.
Remark A.2.
Note that in both cases, since $\mathop{}\!\mathrm{d}\tilde{\mu}(\mathbf{x})=\eta(\mathbf{x})\mathop{}\!\mathrm{d}\mu(\mathbf{x})$ and $\mathop{}\!\mathrm{d}\tilde{\nu}(\mathbf{y})=\xi(\mathbf{y})\mathop{}\!\mathrm{d}\nu(\mathbf{y})$, the equality of mass of $\tilde{\mu}$ and $\tilde{\nu}$ yields $\mathbb{E}_{X\sim\mu}[\eta(X)]=\mathbb{E}_{Y\sim\nu}[\xi(Y)]\,.$
Step 2: Optimality in the balanced problem for the linear case. We now prove point 1, stating that if $\pi^{\star}_{\varepsilon,\tau}$ solves problem ULEOT between $\mu$ and $\nu$, then it solves problem (LEOT) between $\tilde{\mu}$ and $\tilde{\nu}$ for the same entropic regularization strength $\varepsilon$. It also follows from duality. We remind that that thanks to (Séjourné et al., 2023a, Prop. 2)
$\frac{\mathop{}\!\mathrm{d}\pi^{\star}_{\varepsilon,\tau}}{\mathop{}\!\mathrm{d}(\mu\otimes\nu)}(\mathbf{x},\mathbf{y})=\exp\left(\frac{f^{\star}(\mathbf{x})+g^{\star}(\mathbf{y})-c(\mathbf{x},\mathbf{y})}{\varepsilon}\right)$ (10)
with $f^{\star}\in\mathcal{C}(\mathcal{X})$ and $g^{\star}\in C(\mathcal{Y})$. Moreover, by (Nutz, , Theorem 4.2), such a decomposition is equivalent to the optimality in problem (LEOT). Therefore, $\pi^{\star}_{\varepsilon,\tau}$ is solves problem LEOT between its marginals $\tilde{\mu}$ and $\tilde{\nu}$, i.e.
$\pi^{\star}_{\varepsilon,\tau}=\operatorname*{arg\,min}_{\pi\in\Pi(\tilde{\mu},\tilde{\nu})}\int_{\mathcal{X}\times\mathcal{Y}}c(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})+\varepsilon\mathrm{KL}(\pi|\mu\otimes\nu)\,.$ (11)
Step 3: Optimality in the balanced problem for the quadratic case. We now prove point 2, stating that, provided that the costs $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ are conditionally positive (or conditionally negative), if $\pi^{\star}_{\varepsilon,\tau}$ solves problem (UQEOT) between $\mu$ and $\nu$, then it actually solves problem (QEOT) between $\tilde{\mu}$ and $\tilde{\nu}$ for the entropic regularization strength $\varepsilon^{\prime}=m(\pi^{\star}_{\varepsilon,\tau})\,\varepsilon$. We first define the functional:
$\begin{split}&F:(\gamma,\pi)\in\mathcal{M}^{+}(\mathcal{X}\times\mathcal{Y})^{2}\mapsto\int_{(\mathcal{X}\times\mathcal{Y})^{2}}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\gamma(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\,\\ &+\varepsilon\mathrm{KL}(\pi\otimes\gamma|(\mu\otimes\nu)^{2})+\lambda_{1}\mathrm{KL}(\pi_{1}\otimes\gamma_{1}|\mu\times\mu)+\lambda_{2}\mathrm{KL}^{\otimes}(\pi_{2}\otimes\gamma_{2}|\nu\times\nu),\end{split}$ (12)
s.t. $\pi^{\star}_{\varepsilon,\tau}\in\operatorname*{arg\,min}_{\pi\in\mathcal{M}(\mathcal{X}\times\mathcal{Y})}F(\pi,\pi)$. Then, if we define the linearized cost
$c^{\star}_{\varepsilon,\tau}:(\mathbf{x},\mathbf{y})\in\mathcal{X}\times\mathcal{Y}\mapsto\int_{\mathcal{X}\times\mathcal{Y}}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}\mathop{}\!\mathrm{d}\pi^{\star}_{\varepsilon,\tau}(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\,,$ (13)
we get from (Séjourné et al., 2023a, Proposition 9), that $\pi^{\star}_{\varepsilon,\tau}$ solves:
$\begin{split}\pi^{\star}_{\varepsilon,\tau}\in\operatorname*{arg\,min}_{\pi\in\mathcal{M}(\mathcal{\times}\mathcal{Y})}&\int_{\mathcal{X}\times\mathcal{Y}}c^{\star}_{\varepsilon,\tau}(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})+\varepsilon m(\pi^{\star}_{\varepsilon,\tau})\mathrm{KL}(\pi|\mu\otimes\nu)\\ &+\lambda_{1}m(\pi^{\star}_{\varepsilon,\tau})\mathrm{KL}(\pi_{1}|\mu)+\lambda_{2}m(\pi^{\star}_{\varepsilon,\tau})\mathrm{KL}(\pi_{2}|\nu)\,.\end{split}$ (14)
Therefore, $\pi^{\star}_{\varepsilon,\tau}$ solves problem ULEOT between $\mu$ and $\nu$ for a new cost $c^{\star}_{\varepsilon,\tau}$, and the regularization strength $\varepsilon^{\prime}=\varepsilon m(\pi^{\star}_{\varepsilon,\tau})$. We seek to apply point 1, to get that $\pi^{\star}_{\varepsilon,\tau}$ solves problem LEOT between $\tilde{\mu}=p_{1}\sharp\pi^{\star}_{\varepsilon,\tau}$ and $\tilde{\nu}=p_{2}\sharp\pi^{\star}_{\varepsilon,\tau}$ for the same entropic regularization strength $\varepsilon^{\prime}$. To that end, we first verify that $c^{\star}_{\varepsilon,\tau}$ is continuous. Since $c_{\mathcal{X}}$ and $c_{\mathcal{Y}}$ are continuous on $\mathcal{X}\times\mathcal{X}$ and $\mathcal{Y}\times\mathcal{Y}$, the function:
$(\mathbf{x},\mathbf{y},\mathbf{x}^{\prime},\mathbf{y}^{\prime})\in(\mathcal{X}\times\mathcal{Y})^{2}\mapsto|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|^{2}$
is continuous. Therefore, it is bounded since $(\mathcal{X}\times\mathcal{Y})^{2}$ is compact as a product of compact sets. Then, since $\pi^{\star}_{\varepsilon,\tau}$ has finite mass, Lebesgue’s dominated convergence yields the continuity of $c^{\star}_{\varepsilon,\tau}$. We then to apply point 1 and get:
$\pi^{\star}_{\varepsilon,\tau}\in\operatorname*{arg\,min}_{\pi\in\Pi(\tilde{\mu},\tilde{\nu})}\int_{\mathcal{X}\times\mathcal{Y}}c^{\star}_{\varepsilon,\tau}(\mathbf{x},\mathbf{y})\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})+\varepsilon^{\prime}\mathrm{KL}(\pi|\mu\otimes\nu)$ (15)
Since the costs are conditionally positive (or conditionally negative) kernels, (15) finally yields the desired result by applying (Séjourné et al., 2023b, Theorem 3):
$\pi^{\star}_{\varepsilon,\tau}\in\operatorname*{arg\,min}_{\pi\in\Pi(\tilde{\mu},\tilde{\nu})}\int_{\mathcal{X}\times\mathcal{Y}}|c_{\mathcal{X}}(\mathbf{x},\mathbf{x}^{\prime})-c_{\mathcal{Y}}(\mathbf{y},\mathbf{y}^{\prime})|\mathop{}\!\mathrm{d}\pi(\mathbf{x}^{\prime},\mathbf{y}^{\prime})\mathop{}\!\mathrm{d}\pi(\mathbf{x},\mathbf{y})+\varepsilon^{\prime}\mathrm{KL}(\pi|\mu\otimes\nu)$ (16)